how many g of gold should a coin of 33% gold be if when combined with a 5 g pure gold necklace it forms a metal that is 70% gold ?
A thirteen L solution that was nine percent vinegar was mixed with four L
Solution that was sixty-eight percent vinegar. Find the concentration of vinegar
Jason mixed 8mL of a salt solution with 10 mL of a 47% salt solution to make a 45 % salt solution . Find the percent salt concentration of the first solution.

Answer by stanbon(75887)

how many g of gold should a coin of 33% gold be if when combined with a 5 g pure gold necklace it forms a metal that is 70% gold ?
-------
gold + gold = gold
0.33x + 1*5 = 0.70(x+5)
33x + 100*5 = 70x + 70*5
37x = 30*5
x = 4.05 grams (amt. of 33% needed)
--------------------------------------
A thirteen L solution that was nine percent vinegar was mixed with four L
Solution that was sixty-eight percent vinegar. Find the percent salt concentration of the first solution.
-----
vinegar + vinegar = vinegar
0.09*13 + 0.68*4 = x(13+4)
9*13 + 68*4 = 100x*17
----
389 = 1700x
x = 0.2288 = 22.88% (% vinegar in solution)
--------------------------------
Jason mixed 8mL of a ?% salt solution with 10 mL of a 47% salt solution to make a 45 % salt solution .
---
salt + salt = salt
0.0?*8 + 0.47*10 = 0.45(8+10)
-----------
?*8 + 47*10 = 45*8 + 45*10
etc.
Cheers,
Stan H.
---------------

--------------------------